OpenStudy (lexiip):

GEOMETRY EMERGENCY!!! fan+medal+testimonial Prove the Converse of the Pythagorean Theorem using similar triangles.

OpenStudy (lexiip):

@nincompoop @confluxepic @jigglypuff314 @sleepyjess

RhondaSommer (rhondasommer):

I can help! :D

RhondaSommer (rhondasommer):

welcome to os :D

OpenStudy (lexiip):

Thank goodness

RhondaSommer (rhondasommer):

XD So Pythagorean Theorem = a^2 +b^2 = c^2 correct? I have to guide you to the answer so bear with me.

OpenStudy (lexiip):

right

OpenStudy (lexiip):

this is the full question: Prove the Converse of the Pythagorean Theorem using similar triangles. The Converse of the Pythagorean Theorem states that when the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse, the triangle is a right triangle. Be sure to create and name the appropriate geometric figures.

RhondaSommer (rhondasommer):

If the square of one side of a triangle is equal to the sum of the squares of the other two sides, then the triangle is a right triangle so if we have... is it a right triangle?

OpenStudy (lexiip):

no

RhondaSommer (rhondasommer):

okay now show me how you got to that answer using the Pythagorean theorem

OpenStudy (lexiip):

ugh idk I'm so confused

OpenStudy (lexiip):

I've been trying to do this for hours

RhondaSommer (rhondasommer):

so \[a^2+b^2=c^2\] c^2 = hypotenuse. so if it was a right triangle it would be \[3^2+5^2=34^2\] 9+24=1156 34=1156 they arent equal, so it cant be a right triangle

OpenStudy (lexiip):

Oh that was actually what I was doing but it seemed wrong to me

RhondaSommer (rhondasommer):

lets try another triangle and see if you can do it again :D 

OpenStudy (lexiip):

a2+b2=c2 9 + 16 = 25 25 = 2

OpenStudy (lexiip):

25 = 25 sorry

OpenStudy (lexiip):

so it is a right triangle

RhondaSommer (rhondasommer):

so is it a right triangle or not?

OpenStudy (lexiip):

yes

RhondaSommer (rhondasommer):

goodjob. you did it/

RhondaSommer (rhondasommer):

youve proven the theorem.

OpenStudy (lexiip):

so what would be my final answer?

RhondaSommer (rhondasommer):

yu put the examples I gave you down, and then you put down how you solved them, and thats your answer.

OpenStudy (lexiip):

thank you so much :) I'll write a testimonial as soon as I finish the test

OpenStudy (zzr0ck3r):

you did not prove the theorem. You cant prove by example. You need to show its true for the general case.

OpenStudy (zzr0ck3r):

@Rhonasommer That is like if I asked you to prove the square of every number is even. Then you say 2^2 is even and 4^2 is even so it must be true. But of course it is not true.

RhondaSommer (rhondasommer):

"Prove the Converse of the Pythagorean Theorem using similar triangles."

RhondaSommer (rhondasommer):

^ what I did.

OpenStudy (zzr0ck3r):

correct, you did not do that

RhondaSommer (rhondasommer):

o.0 alright. I hand off the baton to you. Good luck! :D

OpenStudy (zzr0ck3r):

no you did not, you proved it for a few example.

OpenStudy (zzr0ck3r):

you need to prove it FOR ALL example.

OpenStudy (zzr0ck3r):

Read the link I posted.

RhondaSommer (rhondasommer):

I handed off the baton :D you explain it. :D

OpenStudy (zzr0ck3r):

The link explains it...

RhondaSommer (rhondasommer):

but we arent supposed to give links. we are supposed to explain it.

OpenStudy (zzr0ck3r):

The link explains it....You can post all the links you want.